OpenStudy (safia21):

Write the standard form of the equation of the line through the given point. through (-3,2) and (0,-1)

OpenStudy (mimi_x3):

m = -1 y = mx + b 2 = -1(-3)+b 2 = 3 + b b = -1 y = -x -1